OpenStudy (anonymous):

A speeding ticket costs $40 plus $5 for each mile per hour over the speed limit. Lynne mentions that she was given a ticket for $75. What is a valid conclusion using deductive reasoning?

OpenStudy (anonymous):

Lynne was driving 7 miles per hour over the speed limit.? Correct.... @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good y = cost x = her speed (over the speed limit) y = 5x + 40 75 = 5x+40 75-40 = 5x 35 = 5x 5x = 35 x = 35/5 x = 7 So you're correct

OpenStudy (anonymous):

thanks, I swear, you're the best tutor on here :)

jimthompson5910 (jim_thompson5910):

yw